OpenStudy (anonymous):

The sequence below shows the number of rocks that Megan has in her rock collection at the end of each week. 11, 16, 21, 26..... Assuming the pattern continues, create an explicit function could be used to find the total number of rocks in her collection at the end of week w? Write your function f(w) in terms of w. Do not use spaces.

OpenStudy (anonymous):

the number of rocks increases 5 each week so after w weeks, he will have f(w) = 11+5w rocks. Let's try: week 0: f(0) = 11 rocks, week 1: f(1)=11+5=16...week 6: f(6)=11+30=41